OpenStudy (anonymous):

Rose bought a new hat when she was on vacation. The following formula describes the percent sales tax Rose paid b=t-c/c where t= total price Rose paid for the hat c= cost of hat b= Sales tax Rewrite the formula to determine the total cost Rose paid for the hat.

OpenStudy (ranga):

Multiply both sides of the equation by c. Gather all the c terms to the left and non-c terms to the right of the = sign. See what division needs to be done to both sides to isolate c on the left side.

OpenStudy (anonymous):

How, would I do that though?

OpenStudy (ranga):

b = (t - c) / c Multiply both sides by c bc = t - c Add c to both sides bc + c = t Factor c out on the left c(b + 1) = t Divide both sides by (b + 1) c = t / (b + 1)

OpenStudy (anonymous):

Okay I get it now thank you!

OpenStudy (ranga):

you are welcome.